The cost of 4 boxes of small paper clips and 9 boxes of large paper clips is $15.75. Eight boxes of small clips and 7 boxes of large ones cost $16.65. Find the cost of a box of each size of paper clips.
I tried to solve this word problem, but with no success. I got that small clips cost -$29.48 and large ones cost $14.85. As you can tell, these answers don't even make sense, and I have no idea what I did wrong! If you could help me that would be great! I have a test on this material tomorrow and I am panicking! :(
Thank You in advance!!!

Answer by scott8148(6628)

"cost of 4 boxes of small paper clips and 9 boxes of large paper clips is $15.75" ___ 4s + 9L = 15.75 ___ 8s + 18L = 31.50

"Eight boxes of small clips and 7 boxes of large ones cost $16.65" ___ 8s + 7L = 16.65

subtracting equations ___ 11L = 14.85 ___ L = 1.35

substituting ___ 4s + 9(1.35) = 15.75 ___ 4s = 3.60 ___ s = .90


the "trick" is to get one set of coefficients to have the same value by multiplying (sometimes both equations)
___ then "cancel" the variable by subtracting